OpenStudy (anonymous):

Please help, I am completely lost. the explanation in the lesson didn't do anything for me. Find the derivative of f(x) = 4x + 7 at x = 5.

OpenStudy (anonymous):

derivative(f(x))=(f(x+h)-f(x))/h as h tends to 0

OpenStudy (anonymous):

in this case, the derivative is (f(5+h)-f(5))/h as h tends to zero.

OpenStudy (anonymous):

i am not really sure what to do with this info tbh :(

OpenStudy (anonymous):

plug in for f. what is f(5+h)? what is f(5)?

OpenStudy (anonymous):

in this case, since the derivative is the slope of the tangent line at a given x value, f'(x)=5.

OpenStudy (anonymous):

you can verify this by calculating it manually. f(5+h)=4*(5+h)+7 f(5)=4*5+7

OpenStudy (anonymous):

f(5)=27

OpenStudy (anonymous):

but what do i do after that?

OpenStudy (anonymous):

calculate f(5+h). subtract f(5) from f(5+h).

OpenStudy (anonymous):

okay give me a moment

OpenStudy (anonymous):

so i solve for h?

OpenStudy (anonymous):

nope. just calculate f(5+h)-f(5) in terms of h.

OpenStudy (anonymous):

i feel retarded right now...

OpenStudy (anonymous):

would you mind working out the problem set by step so that I can use it as reference or link me to a good video on the subject because i have no idea what im doing

OpenStudy (anonymous):

sure. f'(x)=limit of (f(x+h)-f(x))/h, as h->0 f'(5)=(f(5+h)-f(5))/h f(5+h)=4*(5+h)+7=4h+27 f(5)=4*5+7=27 f(5+h)-f(5)=(4h+27)-(27)=4h (f(5+h)-f(5))/h=4h/h=4. Thus the derivative is 4.

OpenStudy (anonymous):

have you studied limits yet?

OpenStudy (anonymous):

i read the lesson in the textbook but i didnt really absorb anything from it

OpenStudy (anonymous):

i get the concepts though

OpenStudy (anonymous):

i just dont really work well without a teacher

OpenStudy (anonymous):

i dont really understand what happens between the last two lines of your explanation f(5+h)-f(5)=(4h+27)-(27)=4h (f(5+h)-f(5))/h=4h/h=4

OpenStudy (anonymous):

i understand the first line but as soon as i see the next line it is a mystery to me

OpenStudy (anonymous):

which part is unclear, (f(5+h)-f(5))=4h, (f(5+h)-f(5))/h=4h/h, or 4h/h=4?

OpenStudy (anonymous):

i think something just clicked in my brain

OpenStudy (anonymous):

can i do one and you check it for me?

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

something came up IRL so i will be back in around 20 minutes if that is okay with you

OpenStudy (anonymous):

Thank you for your help even if you cant wait for me.